a track is in the shape of a rectangle and two semicircles . the ractangle is 4 yards by 10 yards to the nearest yard what is the perimeter if the rack
Answer by ankor@dixie-net.com(22740)

A track is in the shape of a rectangle and two semicircles. The rectangle is 4 yards by 10 yards to the nearest yard what is the perimeter if the track
:
The diameter of the end semicircles = 4 yd
The circumference of the 2 semicircle = pi*4 = 12.566
:
Perimeter = 2(10) + 12.566 = 32.566 yds ~ 33 yds